OpenStudy (axel.caballero):

If f(x) = log3 (x + 1), what is f-1(2) ?

OpenStudy (axel.caballero):

@mathtutoring22 @mathstudent55 Can you help ?

OpenStudy (anonymous):

f^-1(2) = 3^x -1 = (3^2) - 1 = 9-1 = 8

OpenStudy (axel.caballero):

Thank you.